# algorithmic

### v1

@@ Abstract
In this paper, we study the problem of efficient online reinforcement learning in the infinite horizon setting when there is an offline dataset to start with. We assume that the offline dataset is generated by an expert but with unknown level of competence, i.e., it is not perfect and not necessarily using the optimal policy. We show that if the learning agent models the behavioral policy (parameterized by a competence parameter) used by the expert, it can do substantially better in terms of minimizing cumulative regret, than if it doesn’t do that. We establish an upper bound on regret of the exact informed PSRL algorithm that scales as $\tilde{O}(\sqrt{T})$. This requires a novel prior-dependent regret analysis of Bayesian online learning algorithms for the infinite horizon setting. We then propose an approximate Informed RLSVI algorithm that we can interpret as performing imitation learning with the offline dataset, and then performing online learning.
@@ 1 Introduction
Related Work. Recently, offline reinforcement learning, where offline datasets are incorporated (Levine et al., 2020), has attracted some interest due to the widespread practice of pre-training with offline datasets in large language models (Brown et al., 2020; Thoppilan et al., 2022; Hoffmann et al., 2022). One fundamental challenge in this line of research is the distribution shift problem, where the policy learnt from offline data performs poorly in the real world. To address this problem, Jin et al. (2021); Rashidinejad et al. (2021); Xie et al. (2021a) adopted the pessimistic approach, which can be overly conservative in practice. In Uehara and Sun (2021); Rashidinejad et al. (2021); Xie et al. (2021a); Agarwal and Zhang (2022), the authors identified sufficient conditions on the dataset that guarantees a certain level of performance for offline RL algorithms. However, determining whether the dataset meets these conditions can be a complicated problem itself, hence making these results impractical (Kumar et al., 2020; Nair et al., 2020; Argenson and Dulac-Arnold, 2020; Levine et al., 2020; Kostrikov et al., 2021; Wagenmaker and Pacchiano, 2022).
@@ 2 Preliminaries
For a communicating MDP $\mathcal{M}(\theta)$, we use $\bar{r}^{*}(\theta),v(\cdot;\theta)$, and $q(\cdot,\cdot;\theta)$ to denote their corresponding $\bar{r}^{*},v^{*},$ and $q^{*}$ functions, as defined in (2), (3) and (5) respectively. We also make Assumption 2.2, which is the same as that in Ouyang et al. (2017).
@@
Then, we consider a general model wherein the expert only knows a noisy version of the optimal $q$-function. The expert uses the policy (8), except that $q^{*}$ is replaced by $\tilde{q}$, the expert’s perceived state-action value function. The vector $\tilde{q}$ follows Gaussian distribution with mean $q^{*}$ and covariance $\mathbb{I}/\lambda^{2}$, where $\lambda>0$ is the knowledgeability parameter. The two parameters $(\beta,\lambda)$ together will be referred to as the competence of the expert. In this case, we denote the expert’s (stationary) policy as $\pi^{\beta,\lambda}$. Such a model was first introduced in Hao et al. (2023).
The learning agent knows the form of the expert’s policy and the knowledgeability parameter $\lambda$ but may not know the deliberateness parameter $\beta$. When $\beta$ is unknown to the learning agent, we will assume an independent exponential prior for the sake of analytical simplicity, $f_{2}(\beta(s))=\lambda_{2}\exp(-\lambda_{2}\beta(s))$ over $\beta(s)$ where $\lambda_{2}>0$ is the same for all $s$.
@@ 3 The Infinite-horizon Informed PSRL Algorithm for Average MDPs
In the inf-iPSRL algorithm we assume the availability of an planning oracle, called MDPSolve, to output an optimal policy for the MDP $\mathcal{M}(\theta)$ given input $\theta$: We assume that MDPSolve is a deterministic mapping from the space of parameters $\varTheta$ to the space of deterministic stationary Markov strategies $\Pi$ with the following property: $\pi=\texttt{MDPSolve}(\theta)$ satisfies $\pi(s)\in\operatorname*{arg\,max}_{a\in\mathcal{A}}q(s,a;\theta)$ for all $s\in\mathcal{S}$. Note that various dynamic programming algorithms such as value iteration, policy iteration, etc. can be used in MDPSolve.
@@ 4.1 The Informed RLSVI Algorithm for average-reward MDPs
The inf-iPSRL algorithm introduced in the previous section assumes that the posterior updates can be done exactly. In practice, the posterior update in (9) is challenging due to the loss of conjugacy while using the Bayes rule. Thus, we must find a computationally efficient way to do approximate posterior updates (and obtain samples from it) to enable practical implementation. Hence, we propose a novel approach based on Bayesian bootstrapping to obtain approximate posterior samples. The key idea is to perturb the loss function for the maximum a posterior (MAP) estimate and use the point estimate as a surrogate for the exact posterior sample.
Note that in the ensuing, we regard $\beta$ as also unknown to the learning agent (and $\lambda=\infty$ for simplicity). Thus, the learning agent must form a belief over both $\theta$ and $\beta$ via a joint posterior distribution conditioned on the offline dataset $\mathcal{D}_{0}$ and the online data at time $t$, $\mathcal{H}_{t}$. We denote the prior pdf over $\theta$ by $f(\cdot)$ and prior pdf over $\beta$ by $f_{2}(\cdot)$.
Consider the average-reward Bellman optimality eq.:
$\bar{r}^{*}+q^{*}(s,a;\theta)=r(s,a)+\mathbb{E}[\max_{b}q^{*}(s^{\prime},b;\theta)].$
For the sake of compact notation, we denote $q^{*}(s,a;\theta)$ as $q^{\theta}_{t}(s,a)$ in this section. Now, consider the offline dataset,
$\mathcal{D}_{0}=\{(s_{k},a_{k},\check{s}_{k},r_{k})_{k=0}^{N-1}\}$
@@
and let $\theta$ denote the parameterization of the relative q-value function. We introduce the temporal difference error ${\mathcal{E}}_{k}$ (parameterized by a given $q^{\theta}$),
$\mathcal{E}_{k}(q^{\theta}):=\left(r_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right),$
and define recursively,
$\bar{r}_{k+1}:=(1-\eta_{k})\bar{r}_{k}+\eta_{k}r_{k},~{}~{}\bar{r}_{0}:=r_{0}.$
We use this to construct a parameterized offline dataset,
$\mathcal{D}_{0}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=0}^{N-1}\}.$
A parameterized online dataset at time $t$, $\mathcal{H}_{t}(q^{\theta})$ can be similarly defined. To ease notation, we will regard the time $t$ during the online phase as the $(N+t)$ time. Thus,
$\mathcal{H}_{t}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=N}^{N+t}\}\},$
the online dataset available at time $t$.
Note that $q^{\theta}$ is to be regarded as a parameter. Now, at time $t$, we would like to obtain a MAP estimate for $(\theta,\beta)$ by solving the following:
$\begin{split}\textbf{MAP:}~{}~{}~{}\arg\max_{\theta,\beta}~{}&\log P(\mathcal{H}_{t}(q^{\theta})|\mathcal{D}_{0}(q^{\theta}),\theta,\beta)+\log P(\mathcal{D}_{0}(q^{\theta})|\theta,\beta)+\log f(\theta)+\log f_{2}(\beta).\end{split}$ (21)
Denote a perturbed version of the $q^{\theta}$-parameterized offline dataset by
$\tilde{\mathcal{D}}_{0}(q^{\theta})=\{(s_{k},\tilde{a}_{k},\check{s}_{k},\bar{r}_{k},\tilde{\mathcal{E}}_{k})_{k=0}^{N-1}\}$
where we add the following random perturbations: (i) actions have perturbation $w_{l}^{h}\sim\exp(1)$, (ii) rewards have perturbations $z_{k}\sim\mathcal{N}(0,\sigma^{2})$, and (iii) the prior $\tilde{\theta}\sim\mathcal{N}(0,\Sigma_{0})$.
Note that the first and second terms involving $\mathcal{H}_{t}$ and $\mathcal{D}_{0}$ in (21) are independent of $\beta$ when conditioned on the actions. Thus, we have a sum of log-likelihood of TD error, transition and action as follows:
$\begin{split}&\quad~{}\log P(\tilde{\mathcal{D}}_{0}(q^{\theta})|q^{\theta})\\ &=\sum_{k=0}^{N-1}\left(\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log P(\check{s}_{k}|a_{k},s_{k},q^{\theta})+\log P(a_{k}|s_{k},q^{\theta})\right)\\ &\leq\sum_{k=0}^{N-1}\left(\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log\pi^{\beta}(a_{k}|s_{k},q^{\theta})\right).\end{split}$
By ignoring the log-likelihood of the transition term (akin to optimizing an upper bound on the negative loss function), we are actually being optimistic.
For the terms in the upper bound above, under the random perturbations assumed above, we have
$\begin{split}&\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})=-\frac{1}{2}(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k})^{2}+\text{constant}\end{split}$
and
$\begin{split}\log\pi^{\beta}(a_{k}|s_{k},q^{\theta}_{t})&=w_{k}\left(\beta q^{\theta}(s_{k},a_{k})-\log\sum_{b}\exp\left(\beta q^{\theta}(s_{k},b)\right)\right).\end{split}$
Now, denote a perturbed version of the $q^{\theta}$-parameterized online dataset by
~$\tilde{\mathcal{H}}_{t}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\tilde{\mathcal{E}}_{k})_{k=N:N+t}\},$
and thus similar to before, we have
$\begin{split}&\quad~{}\log P(\tilde{\mathcal{H}}_{t}(q^{\theta})|\tilde{\mathcal{D}}_{0}(q^{\theta}),q^{\theta})\\ &=\sum_{k=N}^{N+t}\left(\log P(\tilde{\mathcal{E}}_{k}(q^{\theta})|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log P(\check{s}_{k}|a_{k},s_{k},q^{\theta})\right),\\ &\leq\sum_{k=N}^{N+t}\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta}),\\ \end{split}$
where we again ignored the transition term to obtain an optimistic upper bound.
Given the random perturbations above, we have
@@
$\begin{split}&\quad~{}\log P(\tilde{\mathcal{E}}_{k}(q^{\theta})|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})\\ &=-\frac{1}{2}\left(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right)^{2}+\text{constant}.\end{split}$
The prior over $\beta$, $f_{2}(\beta)$ is assumed to be an exponential pdf $\lambda_{2}\exp(-\lambda_{2}\beta),\beta\geq 0$, while that over $\theta$ is assumed Gaussian. Thus, putting it all together, we get the following optimistic loss function (to minimize over $\theta$ and $\beta$),
@@
~$\begin{split}\tilde{\mathcal{L}}(\theta,\beta)&=\frac{1}{2\sigma^{2}}\sum_{k=0}^{N+T-1}\left(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right)^{2}\\ &-\sum_{t=0}^{T-1}w_{k}\left(\beta q^{\theta}(s_{t},a_{t})-\log\sum_{b}\exp\left(\beta q^{\theta}(s_{t},b)\right)\right)+\frac{1}{2}(\theta-\tilde{\theta})^{\top}\Sigma_{0}(\theta-\tilde{\theta})+\lambda_{2}\beta.\end{split}$ (22)
The above loss function is difficult to optimize in general due to the $\max$ operation, and the $Q$-value function in general having a nonlinear form.
@@
Note that the loss function in (22) can be hard to jointly optimize over $\theta$ and $\beta$. In particular, estimates of $\beta$ can be quite noisy when $\beta$ is large, and the near-optimal expert policy only covers the state-action space partially. Thus, we consider other methods of estimating $\beta$ that are more robust, which can then be plugged into the loss function in (22). Specifically, we could simply look at the entropy of the empirical distribution of the actions in the offline dataset. Suppose the empirical distribution of $\{\bar{a}_{0}^{l},\ldots\bar{a}_{k}\}_{l=1}^{L}$ is $\mu_{A}$. Then we use $c_{0}/\mathcal{H}(\mu_{A})$ as an estimation for $\beta$, where $c_{0}>0$ is a hyperparameter. The intuition is that for smaller $\beta$, the offline actions tend to be more uniform and thus the entropy will be large. This is an unsupervised approach and agnostic to specific offline data generation process.
@@
In the loss function in (22), the parameter $\theta$ appears inside the $\max$ operation. Thus, it can be quite difficult to optimize over $\beta$. Since the loss function is typically optimized via an iterative algorithm such as a gradient descent method, a simple and scalable solution that works well in practice is to use the parameter estimate $\theta$ from the previous iteration inside the $\max$ operation, and thus optimize over $\theta$ only in the other terms.
4.2 inf-iRLSVI  bridges Online RL and Imitation Learning
In the previous subsection, we derived inf-iRLSVI, a Bayesian-bootstrapped algorithm. We now present interpretation of the algorithm as bridging online RL (via commonality with the RLSVI algorithm (Osband et al., 2016) and imitation learning, and hence a way for its generalization.
Consider the RLSVI algorithm for online reinforcement learning as introduced in Osband et al. (2019). It draws its inspiration from the posterior sampling principle for online learning, and has excellent cumulative regret performance. The RLSVI algorithm, that uses all of the data available at the end of episode $t$, including any offline dataset involves minimizing the corresponding loss function at each time step:

### v2

@@ Abstract
In this paper, we study the problem of efficient online reinforcement learning in the infinite horizon setting when there is an offline dataset to start with. We assume that the offline dataset is generated by an expert but with unknown level of competence, i.e., it is not perfect and not necessarily using the optimal policy. We show that if the learning agent models the behavioral policy (parameterized by a competence parameter) used by the expert, it can do substantially better in terms of minimizing cumulative regret, than if it doesn’t do that. We establish an upper bound on regret of the exact informed PSRL algorithm that scales as $\tilde{O}(\sqrt{T})$. This requires a novel prior-dependent regret analysis of Bayesian online learning algorithms for the infinite horizon setting. We then propose the Informed RLSVI algorithm to efficiently approximate the iPSRL algorithm.
@@ 1 Introduction
Related Work. Recently, offline reinforcement learning, where offline datasets are incorporated (Levine et al., 2020), has attracted some interest due to the widespread practice of pre-training with offline datasets in large language models (Brown et al., 2020; Thoppilan et al., 2022; Hoffmann et al., 2022). One fundamental challenge in this line of research is the distribution shift problem, where the policy learnt from offline data performs poorly in the real world. To address this problem, Jin et al. (2021); Rashidinejad et al. (2021); Xie et al. (2021a) adopted the pessimistic approach, which can be overly conservative in practice. In Uehara and Sun (2021); Rashidinejad et al. (2021); Xie et al. (2021a); Agarwal and Zhang (2022), the authors identified sufficient conditions on the dataset that guarantees a certain level of performance for offline RL algorithms. However, determining whether the dataset meets these conditions can be a complicated problem itself, hence making these results impractical (Kumar et al., 2020; Nair et al., 2020; Argenson and Dulac-Arnold, 2020; Levine et al., 2020; Kostrikov et al., 2021; Wagenmaker and Pacchiano, 2022). There have also been a few experimental work on leveraging offline data for online learning (Zheng et al., 2023; Feng et al., 2023; Hu et al., 2023).
@@
Contributions.
(i) We introduced the informed PSRL algorithm for infinite horizon MDPs that incorporates prior data and takes deterministic episode lengths as input (compared to Ouyang et al. (2017) where the episode schedule is fixed and dependent on visitation counts). (ii) We develop an optimal policy estimation-error based technique that differs from the existing information-ratio based techniques (Russo and Van Roy, 2016; Hao et al., 2023) for prior-dependent regret analysis. (iii) We provide a novel prior-dependent regret bound for learning infinite horizon MDPs that scales polynomially with all relevant parameters. (iv) We introduced the iRLSVI algorithm to approximate the informed PSRL algorithm. The algorithm can be seen as bridging online learning and imitation learning.
Notations.
We use script letters for sets. For a set $\mathcal{X}$, we use $\Delta(\mathcal{X})$ to represent the set of probability measures on $\mathcal{X}$. $\mathbb{P},\mathbb{E}$ stand for probability and expectation respectively. We use the standard big-$O$ and big-$\tilde{O}$ notations to hide absolute constants and/or logarithmic terms.
@@ 2 Preliminaries
For a communicating MDP $\mathcal{M}(\theta)$, we use $\bar{r}^{*}(\theta),v(\cdot;\theta)$, and $q(\cdot,\cdot;\theta)$ to denote their corresponding $\bar{r}^{*},v,$ and $q$ functions, as defined in (2), (3) and (5) respectively. We also make Assumption 2.2, which is the same as that in Ouyang et al. (2017).
@@
Remark 2.3.
Note that we can replace the “for all $\theta\in\varTheta$” statements in Assumption 2.1 and 2.2 with “for $\mu_{0}$-almost all $\theta\in\varTheta$” (a.k.a. holds for $\theta^{*}$ with probability 1) without loss of generality: If certain $\theta\in\varTheta$ does not satisfy either assumption we can remove it from the parameter set and it will affect neither our algorithm nor the analysis. [^1] Under certain mild measurability assumptions.
@@
Then, we consider a general model wherein the expert only knows a noisy version of the optimal $q$-function. The expert uses the policy (8), except that $q$ is replaced by $\tilde{q}$, the expert’s perceived state-action value function. The vector $\tilde{q}$ follows Gaussian distribution with mean $q$ and covariance $\mathbb{I}/\lambda^{2}$, where $\lambda>0$ is the knowledgeability parameter. The two parameters $(\beta,\lambda)$ together will be referred to as the competence of the expert. In this case, we denote the expert’s (stationary) policy as $\pi^{\beta,\lambda}$. Such a model was first introduced in Hao et al. (2023).
The learning agent knows the form of the expert’s policy and the knowledgeability parameter $\lambda$ but may not know the deliberateness parameter $\beta$. For the sake of simplicity, when $\beta$ is unknown to the learning agent, we assume that $\beta(s)=\beta$ is an exponential random variables with probability density function $f_{2}(x)=\lambda_{2}\exp(-\lambda_{2}x)\bm{1}_{\{x>0\}}$ where $\lambda_{2}>0$.
@@ 3 The Infinite-horizon Informed PSRL Algorithm for Average MDPs
In the inf-iPSRL algorithm we assume the availability of an planning oracle, called MDPSolve, to output an optimal policy for the MDP $\mathcal{M}(\theta)$ given input $\theta$: We assume that MDPSolve is a deterministic mapping from the space of parameters $\varTheta$ to the space of deterministic stationary Markov strategies $\Pi$ with the following property: $\pi=\texttt{MDPSolve}(\theta)$ satisfies $\pi(s)\in\operatorname*{arg\,max}_{a\in\mathcal{A}}q(s,a;\theta)$ for all $s\in\mathcal{S}$. Note that various dynamic programming algorithms such as value iteration, policy iteration, etc. can be used in MDPSolve. A similar MDP solving oracle is also needed for Ouyang et al. (2017).
@@ 3.1 Prior Dependent Bound
Remark 3.6.
When $\varepsilon=\Theta(1)$, the regret bound in Corollary 3.4 has an additional factor of $\sqrt{SA}$ comparing to the $\tilde{O}(\bar{v}\sqrt{SAT})$ regret bound for (uninformed) learning in communicating MDPs given by Agrawal and Jia (2017). However, the advantage of our prior-dependent bound is for small $\varepsilon$. The next section shows that under certain assumptions, $\varepsilon$ can be exponentially small in the amount of prior data. We leave the improvement of the prior-dependent upper bound’s dependency on $S$ and $A$ as future work.
@@ 4.1 The Informed RLSVI Algorithm for Average-reward MDPs
In the previous section, we considered the inf-iPSRL algorithm where the learning agent performs an posterior update at each episode. However, computing exact posterior updates can be challenging due to the loss of conjugacy when applying Bayes rule in (9). In this section, we propose a novel Baysian bootstrapping inspired approach to perform approximate posterior updates, where we use the maximum a posterior (MAP) estimator based on a perturbed loss function as a surrogate for samples from exact posterior distributions.
In this section, we consider a different setting as the previous section: the expert deliberateness parameter $\beta$ is assumed to be unknown to the learning agent. Instead, $\beta(s)=\beta$ is an exponential random variable with probability density function $f_{2}$. At each time in the algorithm, the learning agent form a joint belief over $(\theta,\beta)$ based on all the offline and online data. The prior distribution of $(\theta,\beta)$ is given by the density function $f\times f_{2}$.
To unify notations for offline and online data, define $\check{s}_{k}$ to be $\bar{s}_{k}$ if $k<N$ and $s_{k-N}$ otherwise). Define $\check{a}_{k}$ similarly. To simplify notations, write $q(s,a;\theta)$, defined in (5), as $q^{\theta}_{t}(s,a)$. Write $r_{k}=r(\check{s}_{k},\check{a}_{k})$.
Consider the average-reward Bellman optimality equation
$\bar{r}^{*}+q(s,a;\theta)=r(s,a)+\mathbb{E}_{s^{\prime}\sim\theta(s,a)}\left[\max_{a^{\prime}\in\mathcal{A}}q(s^{\prime},a^{\prime};\theta)\right].$ (21)
Let $\tilde{r}_{k}$ be defined recursively via
$\tilde{r}_{k+1}:=(1-\eta_{k})\tilde{r}_{k}+\eta_{k}r_{k},~{}~{}\tilde{r}_{0}:=r_{0}.$
@@
Define the temporal difference error ${\mathcal{E}}_{k}$ (parameterized by a given $q^{\theta}$),
$\mathcal{E}_{k}(q^{\theta}):=\left(r_{k}+\max_{a^{\prime}\in\mathcal{A}}q^{\theta}(\check{s}_{k}^{\prime},a^{\prime})-q^{\theta}(\check{s}_{k},\check{a}_{k})-\tilde{r}_{k}\right),$
where $\check{s}_{k}^{\prime}=\check{s}_{k+1}$ for all $k$ except when $k=N-1$, in which case $\check{s}_{k}^{\prime}=\bar{s}_{N}$. Here $\tilde{r}_{k}$ is used as a bootstrap approximation for the true optimal average reward $\bar{r}^{*}$. Therefore, conditioning on $(\theta,\check{s}_{k},\check{a}_{k})$, $\mathcal{E}_{k}(q^{\theta})$ is approximately zero-mean.
Now, for the convenience of describing the algorithm, define the parameterized dataset, which contains offline and online data up to (offline/online combined) time $t$:
$\mathcal{H}_{t}(q^{\theta}):=\{(\mathbf{p}_{k})_{k=0}^{t}\}:=\{(\check{s}_{k},\check{a}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=0}^{t}\}\},$ (22)
At time $t$, consider the following simplistic way of viewing the parameterized dataset: (i) Each $\mathbf{p}_{k}$ is a separate data point where $\check{s}_{k}$ is viewed as a given context. (ii) Conditioning on $(\theta,\check{s}_{k},\check{a}_{k})$, the temporal difference error $\mathcal{E}_{k}(q^{\theta})\sim\mathcal{N}(0,\sigma^{2})$. (iii) The actions in the offline part of the dataset contains additional information (compared to the online part) since the actions were generated through policy $\pi^{\beta}(\cdot|\cdot,q^{\theta})$. Otherwise the actions are also given context. In this view, the maximum a posterior (MAP) estimate of the parameters $(\theta,\beta)$ at (online) time $t$ is given by $\arg\min_{\theta,\beta}\mathcal{L}(\theta,\beta)$ where
$\mathcal{L}(\theta,\beta)$ $=-\sum_{k=0}^{N+t-1}\log\tilde{\mathbb{P}}_{\theta}(\mathcal{E}_{k}(q^{\theta})|\check{s}_{k},\check{a}_{k})-\sum_{k=0}^{N-1}\log\pi^{\beta}(\check{a}_{k}|\check{s}_{k};q^{\theta})-\log f(\theta)-\log f_{2}(\beta)$ (23)
@@
$=\underbrace{\frac{1}{2\sigma^{2}}\sum_{k=0}^{N+t-1}\left(r_{k}+\max_{a^{\prime}\in\mathcal{A}}q^{\theta}(\check{s}_{k},a^{\prime})-q^{\theta}(s_{k},a_{k})-\tilde{r}_{k}\right)^{2}}_{\text{Value Iteration Loss}}$ (24)
@@
$\underbrace{-\sum_{k=0}^{N-1}w_{k}\left(\beta q^{\theta}(\bar{s}_{k},\bar{a}_{k})-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\beta q^{\theta}(\bar{s}_{k},a^{\prime})\right)\right)}_{\text{Imitation Loss}}+\underbrace{\frac{1}{2}\theta^{\top}\Sigma_{0}\theta+\lambda_{2}\beta}_{\text{Prior Loss}}+~{}\text{constant}.$ (25)
Now we add the following random perturbations to the loss function: (i) each temporal difference error term is perturbed by $z_{k}\sim\mathcal{N}(0,\sigma^{2})$, (ii) each imitation loss term is weighted by a random $w_{k}\sim\exp(1)$, and (iii) $\theta$ in the prior loss term is perturbed by $\tilde{\theta}\sim\mathcal{N}(0,\Sigma_{0})$.
Combining the above, we get the following randomized loss function,
~$\begin{split}\tilde{\mathcal{L}}(\theta,\beta)&=\underbrace{\frac{1}{2\sigma^{2}}\sum_{k=0}^{N+t-1}\left(r_{k}+z_{k}+\max_{a^{\prime}\in\mathcal{A}}q^{\theta}(\check{s}_{k},a^{\prime})-q^{\theta}(s_{k},a_{k})-\tilde{r}_{k}\right)^{2}}_{\text{Randomized Value Iteration Loss}}\\ &\underbrace{-\sum_{t=0}^{N-1}w_{k}\left(\beta q^{\theta}(s_{t},a_{t})-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\beta q^{\theta}(s_{t},a^{\prime})\right)\right)}_{\text{Randomized Imitation Loss}}+\underbrace{\frac{1}{2}(\theta-\tilde{\theta})^{\top}\Sigma_{0}(\theta-\tilde{\theta})+\lambda_{2}\beta}_{\text{Randomized Prior Loss}}.\end{split}$ (26)
Algorithm 2 inf-iRLSVI
  $t_{1}\leftarrow 0$
  for $k=1$ to $K$ do
     Construct ~$\tilde{\mathcal{L}}(\theta,\beta)$ as described above
     Solve $(\tilde{\theta}^{k},\tilde{\beta}^{k})=\arg\min_{\theta,\beta}\tilde{\mathcal{L}}(\theta,\beta)$
     $\tilde{\pi}^{k}\leftarrow\texttt{MDPSolve}(\tilde{\theta}^{k})$
     $t_{k+1}\leftarrow t_{k}+T_{k}$
     for $t=t_{k}$ to $t_{k+1}-1$ do
        Take action $a_{t}=\tilde{\pi}^{k}(s_{t})$
     end for
  end for
@@
The MAP estimation of $\beta$ can be quite noisy when $\beta$ is large due to the fact that the expert takes non-optimal actions with very small probability. To resolve this, we may consider an entropy-based estimate of $\beta$: The estimate is set to equal $c_{0}/\hat{H}$, where $\hat{H}$ is the conditional entropy of $\hat{a}$ given $\hat{s}$ if $(\hat{s},\hat{a})$ follows the empirical distribution given by visitation counts in the offline data.
@@
Due to the maximum operation inside and the non-linearity of $q$-function with respect to $\theta$, optimizing (26) can be a difficult problem in general. For iterative optimization algorithms such as gradient descent methods, one simple and efficient solution is to replace the maximum operation with the parameter estimate of $\theta$ in the previous iteration and optimize $\theta$ over other terms.
4.2 Informd RLSVI Bridges Online RL and Imitation Learning
In the previous section, the loss function can be seen as sum of a few different terms. In this section, we present a detailed interpretation of these terms to conclude that the algorithm can be seen as bridging the online RL algorithms (Osband et al., 2019) with imitation learning.
In (Osband et al., 2019), the author considered episodic learning of finite horizon MDPs with horizon $H$. The algorithm was inspired by the posterior sampling approach and has great empirical performance. At the end of episode $t$, the RLSVI algorithm considers randomized loss function